OpenStudy (anonymous):

The diameter of a circle has endpoints P(–10, –2) and Q(4, 6). a.Find the center of the circle. b.Find the radius. If your answer is not an integer, express it in radical form. c.Write an equation for the circle.

OpenStudy (lgbasallote):

center equals midpoint of diameter so \[\large (\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2})\]

OpenStudy (lgbasallote):

\[\large C (\frac{-10 + 4}{2} , \frac{-2 + 6}{2})\] \[\large C (\frac{-6}{2}, \frac{4}{2})\] \[\large C(-3, 2)\]

OpenStudy (lgbasallote):

radius would be diameter divided by 2